OpenStudy (danysanchy97):

On a 50-question test, a student answered 5 questions incorrectly. What percent did the student answer correctly? A. 90% B. 45% C. 55% D. 10%

OpenStudy (fortytherapper):

I would set up a proportion \[\frac{ x }{ 100% }= \frac{ 45 }{ 50 }\] Understand how I got that, or would you like me to explain deeper?

OpenStudy (danysanchy97):

yes please

OpenStudy (fortytherapper):

Alright. So you want to find what percentage he got correctly. A test is worth 100%, right. So what part of the 100% did he get correct. That's where this came from: \[\frac{ x }{ 100 }\] Since we put a equal sign, we must match the correct numbers with the correct percentage. They gave you 5 incorrect answers, but you're trying to find correct answers. If there are 50 total questions and he missed 5, then: \[50-5=45\] There were also 50 total questions, creating this: \[\frac{ 45 }{ 50 }\] Now if we combine then x = 45 because both represent correctness and 100 = 50 because both represent a total \[\frac{ x }{ 100 }=\frac{ 45 }{ 50 }\]

OpenStudy (danysanchy97):

oh ok i understand now. thank you

OpenStudy (fortytherapper):

You're welcome. Now you can just cross-multiply and divide and then that will find X